OpenStudy (anonymous):

How do I graph the function, g(x) = 2log(base 3) (x+2) +3 ?

OpenStudy (anonymous):

looks more or less like the graph of y = ln(x) only shifted up two 3 units and to the left two units. if you have to do it for a class you should plot a couple points. choose them so when you add 2 you get a power of three. for example \[x=-1\] \[g(-1)=2log_3(-1+2)+3=2log_3(1)+3=0+3=3\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

\[x=1\] \[g(1)=2log_2(1+2)+3=2log_3(3)+3=2\times 1 +3=5\]

OpenStudy (anonymous):

hope it is clear what y = ln(x) looks like if not google it to get a nice picture. also hope it is clear why \[log(1)=0\] and why \[log_3(3)=1\]

OpenStudy (anonymous):

so you want your input to be a power of 3. pick x = 7 to get input of 9 etc. also typo in my first line. i meant up 3, left 2